Phenylketonuria, a Metabolic disease in humans, is caused by a recessive allele, k. If two heterozygous carriers of the allele marry and plan a family of five children:

a. What is the chance that all their children will be normal?

b. What is the chance that four children will be normal an done affected?

c. What is the chance that at least three children will be normal?

d. What is th echance that the first child will be a normal girl?


1
Expert's answer
2021-11-10T13:39:01-0500

If they are both heterozygous, then one-quarter of their children would have PKU, so the probability that their first child will have PKU is 1/4, and the probability of their being heterozygous and of their first child having PKU is 4/9 × 1/4 = 4/36 = 1/9, which is the answer to the question.